# Extended AdS spacetime without boundaries and entanglement without holography

### v1

@@ 5 Conclusions
Within this setup, we studied the semiclassical entanglement entropy across an entangling surface defined by two antipodal stacks of $N$ coincident membranes. Our main result is that the entropy satisfies a universal area law, reminiscent of the Ryu–Takayanagi formula, but derived purely from bulk gravitational considerations, without invoking any boundary dual theory. In odd spacetime dimensions, each pair of membranes contributes precisely a quarter of its area to the total entropy, reproducing the familiar prefactor found in black hole thermodynamics and holographic entanglement. In even dimensions, a boundary divergence emerges; however, we showed that this divergence can be canceled by tuning the number of membranes to infinity in a controlled way, effectively regularizing the entropy and restoring the area law.
@@
There are several directions that would be worth exploring further. In particular, it remains an open question whether the mechanism responsible for the cancellation of divergences in the even-dimensional case admits a deeper interpretation—arguably in terms of coarse-graining over microscopic degrees of freedom localized at the defects. Another important question is whether this setup can be embedded in, or derived from, a UV-complete theory of gravity. It would be interesting to investigate whether the extended AdS geometry introduced here can be realized, for instance, as the near-horizon limit of a brane configuration in string theory or M-theory.

### v2

@@ 5 Conclusions
Within this setup, we studied the semiclassical entanglement entropy across an entangling surface defined by two antipodal stacks of $N$ coincident membranes. Our main result is that the entropy obeys a universal area law, analogous to the Ryu–Takayanagi formula, yet derived entirely from bulk gravitational dynamics, without invoking any boundary dual description. Although a UV divergence inevitably arises from the near-boundary region, we showed that this divergence can be canceled in a controlled manner by taking the limit of an infinite number of membranes, thereby regularizing the entropy and restoring the area law. This large-$N$ mechanism for canceling short-distance divergences is the main insight of this work.
@@
There are several directions that would be worth exploring further. In particular, it remains an open question whether the large-$N$ mechanism responsible for the cancellation of divergences admits a deeper interpretation—arguably in terms of coarse-graining over microscopic degrees of freedom localized at the defects. Another important question is whether this setup can be embedded in, or derived from, a UV-complete theory of gravity. It would be interesting to investigate whether the extended AdS geometry introduced here can be realized, for instance, as the near-horizon limit of a brane configuration in string theory or M-theory.